OpenStudy (anonymous):

Write the equation of a circle with a center at (0, 8) and a radius of 9. Write the equation of a circle with a center at (-5, -7) and a radius of 3.8. Write the equation of a circle with a center at (7, 3) and a diameter of 24. Write the equation of a circle with a center at (4, 1) and a radius of 6. Write the equation of a circle with a center at (10, 10) and a radius of 12. Write the equation of a circle with a center at (-3, -4) and a radius of 3. Write the equation of a circle with a center at (6, 7) and a diameter of 4. Please explain me how to do these....

OpenStudy (anonymous):

@DanielK16

OpenStudy (anonymous):

Wouldn't be just easier to learn how to do this yourself? :P Circle = a set of points that are all the same distance from one fixed point.  you're basically deriving equation of a circle from the distance formula, or pythagorus. \[r^2 = (x-h)^2+(y-k)^2\] where the r is the radius and (h,k) is the center.

OpenStudy (anonymous):

so the first question will be 9^2=(x-0)^2 + (y-8)^2 ? @iambatman

OpenStudy (jdoe0001):

\(\large (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{green}{ r}}^2\qquad \begin{array}{rllll} ({\color{brown}{ -5}},{\color{blue}{ -7}})\quad {\color{green}{ r}}=3.8\\ ({\color{brown}{ 7}},{\color{blue}{ 3}})\quad {\color{green}{ r}}=24\\ ({\color{brown}{ 4}},{\color{blue}{ 1}})\quad {\color{green}{ r}}=6\\ ({\color{brown}{ 10}},{\color{blue}{ 10}})\quad {\color{green}{ r}}=12\\ ({\color{brown}{ -3}},{\color{blue}{ -4}})\quad {\color{green}{ r}}=3\\ ({\color{brown}{ 6}},{\color{blue}{ 7}})\quad {\color{green}{ r}}=4\\ \end{array}\)

OpenStudy (anonymous):

Got it !! Thank you so much @jdoe0001 @iambatman